Can you please help me with the following equations. I don't know my negative/positives very well and fractions:
2/7b - 9 = 2
3(22V - 1) = 162
-6 (m+2)+ 2) +4 = 36
5 - 40x = 125

Thank you for your time and help on this.

Answer by sarah_adam(201)

2/7b - 9 = 2
2/7b = 2+9 ( when -ve sign goes to other side of equation it cnages to +ve)
2/7b = 11
2 = 11*7b
b = 2/77

3(22V - 1) = 162
(22v - 1) = 162/3 = 54
22v - 1 = 54
22v = 54 +1 (when -ve sign goes to other side of equation it cnages to +ve)
22v = 55
v = 55/22 = 5/2
I didnt understand another bracket ending after 2 when there is no beginning bracket. so iam assuming the problem to be
-6 (m+2)+ 2 +4 = 36
-6(m+2) +6 = 36
-6(m+2)=36-6(when +ve sign goes to other side of the equation it changes to
-ve)
-6(m+2) = 30
-(m+2)= 30/6 = 5
-m - 2 = 5
-m = 5+2 = 7
m = -7

5 - 40x = 125
-40x = 125 -5
-40x = 120
-x = 120/40
-x = 3
x = -3